OpenStudy (anonymous):

y = -3x^2 + 12x + -5. I need to find the x-coordinate of the vertex of this parabola? But I have not been able to figure out how to factor this particular equation. What are the factors? and x Coordinates?

OpenStudy (anonymous):

vertex V=(2,7)

OpenStudy (anonymous):

Mark can I see how you found the x and y coordinates?

OpenStudy (anonymous):

verex x=-b/2a x=-12/2(-3)=12/6=2, put this x value to the equation to get the y of the vertex y = -3x^2 + 12x + -5. y = -3(2)^2 + 12(2) -5. y=7 therfore V(2,7) ans

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw good luck joseph